Conduct a t-test for the null hypothesis against



alternative hypothesis based on 16 random



observations. The sample mean is 110 and the sample standard



deviation is 10. Use α = 0.05.

1
Expert's answer
2022-06-15T14:25:15-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=100"

"H_1:\mu>100"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=15" and the critical value for a right-tailed test is "t_c =1.75305."

The rejection region for this right-tailed test is "R = \{t:t>1.75305\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{110-100}{10/\sqrt{16}}=4"


Since it is observed that "t=4>1.75305=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=15" degrees of freedom, "t=4" is "p=0.00058," and since "p=0.00058<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 100, at the "\alpha = 0.05" significance level.